# Are penalty shootouts better than a coin toss? Evidence from international club football in Europe

### v4

@@ 5.2 Multi-dimensional analysis
Table 2 investigates the effect of strength difference between the teams on the probability of winning a penalty shootout by logistic regressions, similar to Wunderlich et al., (2020). We study only matches with a home and an away team (Table 2.a uses an extended and Table 2.b uses a restricted definition), and control for first-mover advantage. The consideration of the momentum effect reduces the number of observations since drawn matches or matches with a final result of 0-0 are excluded. The model containing comeback teams identified by scoring more goals does not depend on the definition of home venue, as the 20 matches affected by the Covid-19 pandemic and a particular final are all played in one leg (see Section 3), thus, their final result should be a draw if they are finished by a penalty shootout.
In Table 2, no significant coefficient can be found. Nonetheless, the constant is always negative in line with the results of binomial tests in Figure 3, indicating a potential home disadvantage. Albeit insignificant, the coefficient $\beta_{2}$ is always positive, while $\beta_{3}$ is negative/positive when the momentum effect is identified by scoring all goals/the last goal, as expected from the binomial tests. Crucially, the coefficient $\beta_{1}$ of strength difference remains robustly insignificant.
We have carried out analogous logistic regression such that the comeback team is identified by both scoring more goals and scoring the last goal, as well as by controlling for goal difference (home team goals minus away team goals) instead of the trichotomous variable of scoring more goals, analogous to Krumer, (2021). These results are not reported because no estimation contains any significant coefficient, even at the 10% level.
@@
Note: The dotted (dashed) line shows significance at 5% (1%) level.
Last but not least, we have assessed whether team strength could have a significant effect by focusing on the recent seasons. Hence, 25-25 logistic regressions are estimated by controlling for strength difference and first-mover effect with the extended definition of home field: the sample begins in the 20xx/20xx+1 season and ends in 2025. Figure 5 plots the corresponding $p$-values for the coefficients $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$. The winning probability becomes different from a coin toss when all seasons before 2023/24 are ignored—but then the sample size is not more than 64. Furthermore, $\beta_{1}$ is significant only if the last one or two season(s) are considered together with the first half of the 2025/26 season. Even then, both $\beta_{1}$ and $\beta_{2}$ are negative, which indicates rather a statistical anomaly caused by inadequate sample size. Consequently, in contrast to domestic cups, stronger teams are not expected to win more penalty shootouts in UEFA club competitions.
5.3 Discussion
@@
In contrast to all existing studies, stronger teams are found not to win more often, which can be deemed counterintuitive. However, we do not see such a problem after a more detailed analysis. First, Arrondel et al., (2019) and Krumer, (2020) find that teams from a higher-ranked division have a higher probability of winning in a penalty shootout. Our sample is more homogeneous from this perspective: teams from the highest-ranked European leagues do not play in the UEFA Champions League qualification since the 2018/19 season, and at most one team participates in the UEFA Europa League and UEFA (Europa) Conference League qualification. Furthermore, the UEFA Champions League qualification was separated into two paths in the 2009/10 season such that champions from lower-ranked national associations are not allowed to play against non-champions from higher-ranked national associations (Csató,, 2022). Even though the knockout stage does not contain analogous constraints, the previous group stage/league phase can effectively remove weak clubs, guaranteeing that teams playing against each other are quite close in abilities. Second, although Wunderlich et al., (2020) report a significant impact for betting odds in domestic cups, its magnitude is relatively small, even extremely weaker teams have a probability of around 40% to win the penalty shootout. UEFA club competitions are more prestigious and lucrative than national clubs, thus strong teams probably field their best squad. Therefore, if the strength difference between the opposing teams is substantial, the match is less likely to be decided by a penalty shootout in UEFA club competitions than in national cups.
@@
Nevertheless, it is important not to over-interpret our finding that stronger teams do not win in a penalty shootout with a higher probability than their weaker opponents. First, the whole analysis is conditional on matches reaching a penalty shootout. Team strength certainly increases the probability of winning earlier, as shown by Csató, (2024), which implies that shootouts involve more closely matched teams. Second, we have focused on one measure of team strength, Football Club Elo Ratings. Using alternative measures of strength, such as UEFA club coefficients or market values, may change the results.
6 Concluding remarks
Our paper has aimed to predict success in penalty shootouts in top European football during the last 25 years. We have considered several possible factors, such as the shooting order, the field of the match, and psychological momentum, but none of them affects the probability of winning significantly. First in the literature, team strength is measured by Elo ratings instead of the division of the team or betting odds. This has led to a somewhat surprising finding: even though Elo ratings can forecast results in the UEFA Champions League relatively well, they seem to lose their predictive power if the qualification is decided by penalty shootouts.
@@ A.2 Supplementary tables
Tournament phase Qualification Knockout Potential success factor Total Ratio $p$-value Total Ratio $p$-value First-mover 207 52.2% 0.578 61 57.4% 0.306 Home (extended data) 204 46.1% 0.294 51 47.1% 0.780 Home (restricted data) 184 47.3% 0.507 50 48.0% 0.888 Comeback (all goals) 108 49.1% 0.923 39 48.7% 1.000 Comeback (last goal) 176 53.4% 0.407 55 58.2% 0.281 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$. The column Total (Ratio) shows the number (proportion) of penalty shootouts won by the team indicated in the row.
@@
Tournament phase Qualification Knockout Scenario Total Ratio $p$-value Total Ratio $p$-value Elo difference >0>0 205 48.3% 0.675 61 42.6% 0.306 Elo difference >25>25 188 49.5% 0.942 54 40.7% 0.220 Elo difference >50>50 165 49.1% 0.876 48 41.7% 0.312 Elo difference >75>75 145 46.9% 0.507 34 47.1% 0.864 Elo difference >100>100 134 47.0% 0.546 20 50.0% 1.000 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$. The column Total (Ratio) shows the number (proportion) of penalty shootouts won by the favourite team.

### v5

@@ 5.2 Multi-dimensional analysis
Table 2 investigates the effect of strength difference between the teams on the probability of winning a penalty shootout by logistic regressions, similar to Wunderlich et al., (2020) and van Hemert and van der Kamp, (2026). We study only matches with a home and an away team (Table 2.a uses an extended and Table 2.b uses a restricted definition), and control for first-mover advantage. Crucially, no significant coefficient can be found. Nonetheless, the constant is always negative in line with the results of binomial tests in Figure 3, indicating a potential home disadvantage. Albeit insignificant, the coefficient $\beta_{2}$ is always positive, while $\beta_{3}$ is negative/positive when the momentum effect is identified by scoring all goals/the last goal, as expected from the binomial tests. The coefficient $\beta_{1}$ of strength difference also remains robustly insignificant.
As mentioned in Section 4, we have carried out analogous logistic regression such that the comeback team is identified by both scoring more goals and scoring the last goal, as well as by controlling for goal difference (home team goals minus away team goals) instead of the trichotomous variable of scoring more goals, analogous to Krumer, (2021). These results are reported in Table A.5; no estimation contains any significant coefficient, even at the 10% level.
@@
Note: The dotted (dashed) line shows significance at the 5% (1%) level.
Last but not least, we have assessed whether team strength could have a significant effect by focusing on the recent seasons. Hence, 25-25 logistic regressions are estimated by controlling for strength difference and first-mover effect with the extended definition of home field: the sample begins in the 20xx/20xx+1 season and ends in 2025. Figure 5 plots the corresponding $p$-values for the coefficients $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$. The null hypothesis of equal winning probabilities can be rejected only when all seasons before 2023/24 are ignored—but then the sample size is not more than 64. Furthermore, $\beta_{1}$ is significant only if the last one or two season(s) are considered together with the first half of the 2025/26 season. Even then, both $\beta_{1}$ and $\beta_{2}$ are negative, which indicates rather a statistical anomaly caused by inadequate sample size. Consequently, in contrast to domestic cups, stronger teams are not expected to win more penalty shootouts in UEFA club competitions.
6 Discussion and concluding remarks
@@
In contrast to all existing studies, stronger teams are found not to win more often, which can be deemed counter-intuitive. However, we do not see such a problem after a more detailed analysis. First, Arrondel et al., (2019) and Krumer, (2020) find that teams from a higher-ranked division have a higher probability of winning in a penalty shootout. Our sample is more homogeneous from this perspective: teams from the highest-ranked European leagues do not play in the UEFA Champions League qualification since the 2018/19 season, and at most one team participates in the UEFA Europa League and UEFA (Europa) Conference League qualification. Furthermore, the UEFA Champions League qualification was separated into two paths in the 2009/10 season, such that champions from lower-ranked national associations are not allowed to play against non-champions from higher-ranked national associations (Csató,, 2022). Even though the knockout stage does not contain analogous constraints, the previous group stage/league phase can effectively remove weak clubs, guaranteeing that teams playing against each other are quite close in abilities. Second, although Wunderlich et al., (2020) report a significant impact for betting odds in domestic cups, its magnitude is relatively small, even extremely weaker teams have a probability of around 40% to win the penalty shootout. UEFA club competitions are more prestigious and lucrative than national clubs, thus strong teams probably field their best squad. Therefore, if the strength difference between the opposing teams is substantial, the match is less likely to be decided by a penalty shootout in UEFA club competitions than in national cups.
@@
Nevertheless, it is important not to over-interpret our finding that stronger teams do not win in a penalty shootout with a higher probability than their weaker opponents. First, the whole analysis is conditional on matches reaching a penalty shootout. Team strength increases the probability of winning earlier (Csató,, 2024), which implies that shootouts involve more closely matched teams, as verified by recent empirical evidence in UEFA club competitions (Forrest et al.,, 2025). Second, we have focused on one measure of team strength, Football Club Elo Ratings. Using alternative measures of strength, such as UEFA club coefficients or market values, may change the results.
To conclude, the current paper has aimed to predict success in penalty shootouts in UEFA club competitions during the last 25 years. We have considered several possible factors, such as the shooting order, the field of the match, and psychological momentum, but none of them affects the probability of winning significantly. First in the literature, team strength is measured by Elo ratings instead of the division of the team or betting odds. This has led to a somewhat surprising finding: even though Elo ratings can forecast results in the UEFA Champions League relatively well, they seem to lose their predictive power if qualification is decided by penalty shootouts.
@@ A.2 Supplementary tables
Tournament phase Qualification Knockout (including the final) Potential success factor Total Ratio $p$-value Total Ratio $p$-value First-mover 207 52.2% 0.578 61 57.4% 0.306 Home (extended data) 204 46.1% 0.294 51 47.1% 0.780 Home (restricted data) 184 47.3% 0.507 50 48.0% 0.888 Comeback (all goals) 108 49.1% 0.923 39 48.7% 1.000 Comeback (last goal) 176 53.4% 0.407 55 58.2% 0.281 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$. The column Total (Ratio) shows the number (proportion) of penalty shootouts won by the team indicated in the row.
@@
Tournament phase Qualification Knockout (including the final) Scenario Total Ratio $p$-value Total Ratio $p$-value Elo difference >0>0 205 48.3% 0.675 61 42.6% 0.306 Elo difference >25>25 188 49.5% 0.942 54 40.7% 0.220 Elo difference >50>50 165 49.1% 0.876 48 41.7% 0.312 Elo difference >75>75 145 46.9% 0.507 34 47.1% 0.864 Elo difference >100>100 134 47.0% 0.546 20 50.0% 1.000 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$. The column Total (Ratio) shows the number (proportion) of penalty shootouts won by the favourite team.
Table A.5: Logistic regression models for predicting success in penalty shootouts
Robustness check with alternative definitions of the comeback team
a Extended definition of home-away matches
Constant ($\beta_{0}$) -0.113 -0.116 -0.066 -0.068
(0.139) (0.141) (0.348) (0.345)
Strength difference ($\beta_{1}$) 0.009 0.009 -0.096 -0.098
(0.077) (0.076) (0.642) (0.636)
First-mover ($\beta_{2}$) 0.033 0.030 0.030 0.027
(0.134) (0.135) (0.134) (0.135)
Comeback (double, $\beta_{3}$) 0.059 0.036
(0.218) (0.200)
Comeback (goal diff., $\beta_{3}$) 0.037 0.031
(0.140) (0.140)
Observations 232 232 232 232
  • Comeback (double) is 1 (1-1) if the home (away) team scored both more goals and the last goal.
  • Comeback (goal diff.) is the number of goals scored minus the number of goals conceded for the home team.
b Restricted definition of home-away matches
Constant ($\beta_{0}$) -0.158 -0.160 -0.168 -0.168
(0.133) (0.135) (0.327) (0.325)
Strength difference ($\beta_{1}$) 0.023 0.022 0.017 0.013
(0.073) (0.072) (0.606) (0.601)
First-mover ($\beta_{2}$) 0.116 0.114 0.114 0.112
(0.128) (0.129) (0.128) (0.129)
Comeback (double, $\beta_{3}$) 0.055 0.044
(0.199) (0.199)
Comeback (goal diff., $\beta_{3}$) 0.040 0.034
(0.140) (0.139)
Observations 253 253 253 253
  • Comeback (double) is 1 (1-1) if the home (away) team scored both more goals and the last goal.
  • Comeback (goal diff.) is the number of goals scored minus the number of goals conceded for the home team.